OpenStudy (anonymous):

Your friend has $100 when he goes to the fair. He spends $10 to enter the fair and $20 on food. Rides at the fair cost $2.00 per ride. Which function can be used to determine how much money he has left over after x rides?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

(100-30) - 2x= answer

OpenStudy (anonymous):

how

OpenStudy (anonymous):

f(x) = -30x + $100 ?

OpenStudy (anonymous):

because first he starts with $100 and then 20 + 10 = 30 so then 100-30 equal 70 and then they did not say how many rides he went so it it is x times $2.00

OpenStudy (anonymous):

so the equation is 70-2.00x= answer

OpenStudy (anonymous):

wouldnt the 2x be negative

OpenStudy (anonymous):

no

OpenStudy (anonymous):

f(x) = -2x + 70 this is the correct answer but thanks.

OpenStudy (anonymous):

oh ok